OpenStudy (anonymous):

Solve. 3x2 = -15x + 21

OpenStudy (anonymous):

3x^2 = -15x + 21 x^2=-5x+7 x^2-7+5x=0 \[x=\frac{ -5\pm \sqrt{53} }{ 2 }\]

OpenStudy (anonymous):

thank you so much!